As a car ages, its value decreases. The value of a particular car with an original purchase price of $27,200 is modeled by the following function, where c is the value at time t, in years.
C(t)=27,200(1-0.22)^t
A. What is the value of the car when it is 3 years old? The value of the car after 3 years is $
B. What is the total depreciation amount after 5 years? After 5 years, the total depreciation amount is $

Answer by Theo(13342)

the formula for value of a car after t years is c(t) = 27,200 * (1 - .22) ^ t
c is the value of the car.
t is the number of years.
what this formula is saying is that the car loses 22% of its value every year.
when t = 3, the formula becomes c(t) = 27,200 * (1 - .22) ^ 3 = 12,907.8144.
when t = 5, the formula becomes c(t) = 27,200 * (1 - .22) ^ 5 = 7,853.114281.
after 5 years, the total depreciaton amount is 27,200 minus 7,853.114281 = 19,346.88572.